OpenStudy (anonymous):

Studying for a math test can someone step by step show me how to find the x and y int. on y=-5x-4

TheSmartOne (thesmartone):

x-intercept is when y=0

TheSmartOne (thesmartone):

To find the x-intercept y has to equal to 0 So 0=-5x-4 add 4 on both sides 4=-5x x=4/-5

TheSmartOne (thesmartone):

to find the y-intercept x has to equal to 0 y=-5(0)-4 y=-4

TheSmartOne (thesmartone):

So the points are (0,-4) and (4/-5,0)

TheSmartOne (thesmartone):

@Syli

OpenStudy (anonymous):

Ok thank you so much kind person.